OpenStudy (anonymous):

Simplify: (q/(q^2-r^2))/r/(r^2-q^2)

OpenStudy (anonymous):

\[\frac{q }{(q-r)(q+r) }*\frac{ r^2-q^2}{r }\\ \frac{q }{ (q-r)(q+r) }*\frac{( r-q)(r+q )}{r }\\\\ \frac{q }{ (q-r)(q+r) }*\frac{-1( q-r)(q+r )}{r }\\\\ \frac{q }{ (\cancel{q-r)(q+r)} }*\frac{-1\cancel{( q-r)(q+r )}}{r }\\\\ =\frac{-q }{ r } \]

OpenStudy (anonymous):

What if it was addition instead of multiplication?

jhonyy9 (jhonyy9):

when you need divide two fractions just multiplie first by second inversed